Given two displacement A=( 8i+4j-2k)m & B = (4i - 6j + 8k)m gind th magnitufe of the displacement 3A-B


1
Expert's answer
2021-07-19T09:50:33-0400
"\bold{3A-B}=3(8,4,-2)-(4,-6,8)=(20,18,-14)\\|\bold{3A-B}|=\sqrt{20^2+18^2+14^2}=30.3"